A 50.0 kg woman, riding on a 10 kg cart, is moving east at 5.0 m/s. The woman  jumps off the front of the cart and lands on the ground at 7.0 m/s eastward. What is  the final velocity of the cart?


1
Expert's answer
2021-01-12T12:11:24-0500

By Conservation Impulse Law:

"(m_1+m_2)v=m_1v_1+m_2v_2"

where "m_1" is mass of cart, "m_2" is mass of woman; "v" is velocity of cart with woman before jumping, "v_1" and "v_2" are velocities of cart and woman after jumping.

"(50+10)\cdot5=50\cdot7-10v_1"

"v_1=\frac{50\cdot7-60\cdot5}{10}=5 m/s" westward.